OpenStudy (anonymous):

Show that the equation has a root in the given interval. -x^4+3x^2+5=0 [2,3]

OpenStudy (anonymous):

Plug in 2, and then 3. The sign of the function will change. That means it's cut the x-axis at least once, and so there is at least one real root in that interval.

OpenStudy (anonymous):

f(2) = 1 f(3) = -49 Look at the graph between x=1 and x=3, you'll see it shoots straight down.

OpenStudy (anonymous):

Am I plugging it as this: f(2) = (-2)^4 + 3(2)^2 +5

OpenStudy (anonymous):

No. \[f(2)=-(2)^4+3(2)^2+5=-16+12+5=1\]

OpenStudy (anonymous):

You replace x with 2, not -2.

OpenStudy (anonymous):

Same for 3:\[f(3)=-(3)^4+3(3)^2+5=-81+27+5=-49\]

OpenStudy (anonymous):

Okay, Jennifer?

OpenStudy (anonymous):

I have to log off. Just be careful when you sub. numbers in for x...

OpenStudy (anonymous):

Okay, th anks